A 40.0-kg cannonball leaves a 2.00 x 10^3 kg cannon with a velocity of +8.00 x 10^1 m/s. What is the recoil velocity of the cannon? 





1
Expert's answer
2021-02-23T10:04:02-0500

Let,

Mass of cannon be "m_e" = 2000kg

Mass of cannon ball be "m_c" = 40kg

Velocity of cannon ball "v_b" = 80 m/s

Velocity of cannon = "v_c"


Using the conservation of Angular momentum


"Initial\ momentum = Final\ momentum"


"P_i = P_f"


"m_bv_b + m_cv_c = 0"


"v_c = -\dfrac{m_b}{m_c}(v_b)"


="- \dfrac{40}{2000}\times 80"


"v_c = -1.6 m\/s"